Sketch the region of integration, carefully with appropriate labels, and then perform the following integral: (0 to pi/4) ∫ (4secƟ to 0) ∫ r dr dƟ
1
Expert's answer
2020-11-01T17:57:11-0500
"\int_0^{\pi/4}\int_{4\sec(\theta)}^0r\,dr\,d\theta=\int_0^{\pi/4}\bigg[r^2/2\bigg]_{4\sec(\theta)}^0d\theta\\
=\int_0^{\pi/4}-8\sec^2(\theta)d\theta\\
=-8[\tan(\theta)]_0^{\pi/4}=-8"